OpenStudy (anonymous):

A parallelogram whose vertices have coordinates R(1, -1), S(6, 1), T(8, 5), and U(3, 3) has a shorter diagonal of ___ . 5 √(13) √(97)

OpenStudy (anonymous):

@perl can you please help??

OpenStudy (perl):

one moment

OpenStudy (anonymous):

Thank you!!

OpenStudy (perl):

I plotted the points here

OpenStudy (perl):

thats the short diagonal http://prntscr.com/7izzgt

OpenStudy (anonymous):

So would it be 5?

OpenStudy (perl):

we want the distance between (3,3) and (6,1) distance = \( \sqrt{ (6-3)^2 + (1-3)^2} = \sqrt{ 9 + 4} = \sqrt{13}\)

OpenStudy (anonymous):

Ugh. I did it wrong, Thank you for helping me! If I have any more can you help me?

OpenStudy (perl):

ok :)